A particular transformer has fewer turns of wire on the secondary coil than on the primary coil. Which of the following statements is true?

The primary voltage is greater than the secondary voltage
The secondary power is greater than the primary power
The length of the secondary coil is greater than the length of the primary coil
The primary current is greater than the secondary current
Answer #1

Ans: The primary voltage is greater than the secondary voltage
Because:
Vp/Vs= Np/Ns
if, Ns<NP
then Vs<Vp
Answer #2

A particular transformer has fewer turns of wire on the secondary coil than on the primary coil

V1/V2 = N1/N2

GIve that N2 < N1

V1/V2 > 1

V1 > V2

Hence answer is

a) The primary voltage is greater than the secondary voltage
Answer #3

The primary voltage is greater than the secondary voltage
Answer #4

The primary voltage is greater than the secondary voltage